OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2ð). cos 4x - cos 2x = 0

OpenStudy (anonymous):

Let u= 2x \[\cos(2u)=\cos^2(u)-\sin^2(u)=2\cos^2(u)-1\] It's not a quadratic.

OpenStudy (anonymous):

What do you mean it's not quadratic?

OpenStudy (anonymous):

It IS quadratic.

OpenStudy (anonymous):

0 pi/3 2pi/3 4pi/3 5pi/3, pi?

OpenStudy (anonymous):

\[2\cos^2(u)-\cos(u)-1=0\] \[(2\cos(u)+1)(\cos(u)-1)=0\] \[\cos(u)-1=0, \cos(u)=1, u=2x=0,2 \pi\] or \[2\cos(u)+1=0, \cos(u)=-0.5, u=2x=5\pi/8, 7 \pi/8\]

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

What part?

OpenStudy (anonymous):

I'm confused by the u

OpenStudy (anonymous):

I simply called the quantity '2x' 'u' for simplicity.

OpenStudy (anonymous):

I just realized that, I think I got it! Thank you for the help